# Scalable magnetic resonance fingerprinting: Incremental inference of high dimensional elliptical mixtures from large data volumes

### v1

@@ 3.2 Contribution
Herein, we propose to explore this divide-and-conquer strategy by introducing the framework of High Dimensional Mixtures of Elliptical Distributions (HD-MED) to get over the hump of estimating more than $2$ parameters, e.g. $T_{1}$, $T_{2}$, $\delta\!f$, $B_{1}$, CBV, $R$. Considering the reconstruction of 6 parameters makes the associated dictionaries both large in size (an order of $13$ terabyte of signals to considerably represent the tissues heterogeneity) and dimension, making our proposal twofold. Building on HD-MED, a generalization of HD-GMM, we show how they can be used to simultaneously compress and cluster large-scale high-dimensional MRF dictionaries, and more generally any dataset. We derive a new incremental algorithm, based on a principled EM framework, to learn such a model from potentially very large data volumes. We demonstrate the effectiveness of our approach on MRF reconstruction, showing results comparable to the high-dimensional dictionary matching referred to as full-matching in the next sections. This approach allows us to exceed the resolution, size used in current implementations and to reconstruct a larger number of MR parameter maps with improved accuracy, thereby advancing the clinical feasibility of MRF.
@@ 4 High Dimensional Mixtures of Elliptical Distributions
Elliptical distributions represent a family of distributions that contains Gaussian distributions but also heavy-tailed distributions, such as the Student distributions, that are often used as a more robust alternative to the Gaussian family. In this paper, we consider a sub-class of elliptical distributions, which can be expressed as infinite mixtures of Gaussian distributions.
@@ 4.1 Gaussian scale mixture (GSM) distributions
Scale mixtures of multivariate Gaussian distributions are an important subclass of elliptical distributions whose definition is recalled below. Scale mixtures of Gaussians share good properties with Gaussian distributions. They are tractable, lead to tractable inference procedures and provide more robust results, in contrast to Gaussian distributions that usually suffer from sensitivity to outliers.
@@
A continuous random vector $\mathbf{Y}\in\mathbb{R}^{M}$ follows a multivariate elliptical symmetric distribution if its probability density function (pdf) $p(\mathbf{y})$ is of the following form (see Cambanis, Huang and Simons (1981) or Kelker (1970)),
@@
Note that the scale matrix ${\boldsymbol{\Sigma}}$ is not necessarily equal to the covariance matrix, ${\boldsymbol{\Sigma}}$ is proportional to the covariance matrix if the latter exists. The pdf of the multivariate normal distribution is a special case of ED with $g(u)=\exp(-u^{2}/2)$. Another member of the elliptical family is the multivariate Student distribution. This distribution is well-studied in the literature (Kotz and Nadarajah, 2004) and admits a useful representation as a Gaussian scale mixture. Denoting ${\cal N}_{M}(\mathbf{y};{\boldsymbol{\mu}},{\boldsymbol{\Sigma}})$ a $M$-variate Gaussian distribution with mean ${\boldsymbol{\mu}}$ and covariance matrix ${\boldsymbol{\Sigma}}$, a Gaussian scale mixture distribution is a distribution of the following form.
@@
If ${\boldsymbol{\mu}}$ is a $M$-dimensional vector, ${\boldsymbol{\Sigma}}$ is a $M\times M$ positive definite symmetric matrix and $f$ is a pdf of a univariate positive variable $W\in\mathbb{R}^{+}$, then the $M$-dimensional density given by
[…]
@@
In practice, we will consider mixing distribution $f_{\boldsymbol{\theta}}$ that depends on some parameter ${\boldsymbol{\theta}}$ and also write $\mathbf{Y}\sim{\cal E}_{M}\left({\boldsymbol{\mu}},{\boldsymbol{\Sigma}},{\boldsymbol{\theta}}\right)$. As already mentioned, famous GSM distributions include, the multivariate Student distribution (when $f_{\boldsymbol{\theta}}$ is the pdf of a Chi2 variable), the Pearson type VII distribution (when $f_{\boldsymbol{\theta}}$ is a gamma distribution) and the generalized Gaussian (when a power of $W$ follows the gamma distribution). It is straightforward to see that GSM are elliptical distributions. However, not all elliptical distributions can be reduced to scale mixtures. For the previous reason, characterization and a way to represent elliptical distributions as GSM are very valuable. In Gómez-Sánchez-Manzano, Gómez-Villegas and Marín (2006), conditions are given under which elliptical distributions are GSMs. The issue of finding the corresponding mixing distribution is also addressed. An illustration of these results for generalized Gaussian distributions is given by Gomez, Gomez-Villegas and Marin (2008).
@@ 4.2 Mixtures of High Dimensional GSM
Consider a data set of $N$ independent observations $\{\mathbf{y}_{1},\dots,\mathbf{y}_{N}\}\in\mathbb{R}^{M}$ assumed to be i.i.d. realizations of a random vector $\mathbf{Y}\in\mathbb{R}^{M}$. In addition, the data set is assumed to be made of $K$ groups to be discovered. In this work, we consider finite mixtures of GSM, assuming that the underlying subsets are distributed according to different GSM in different proportions. In addition, to handle potentially high-dimensional observations with GSM, we propose a specific parameterization of the scale matrices.
Definition 4.3 (Finite mixture model).
@@
where $\pi_{k}\in[0,1]$ are the mixing weights that sum to one, and $f_{k}$ is the conditional pdf of the $k^{\text{th}}$ mixture component.
To simplify the notation, a finite mixture model where each pdf $f_{k}$ is the pdf of a GSM distribution (3), ${\cal E}({\boldsymbol{\mu}}_{k},{\boldsymbol{\Sigma}}_{k},{\boldsymbol{\theta}}_{k})$, is referred to as a mixture of ED (MED). The number of parameters in MED grows quadratically with the dimension $M$ due to the scale matrices ${\boldsymbol{\Sigma}}_{k}$. For large $M$, this can be problematic for the mixture estimation from a data set. In the Gaussian mixture case, to reduce the number of parameters, Bouveyron, Girard and Schmid (2007) proposed a family of parsimonious Gaussian mixture models, using the eigenvalues decomposition of the covariance matrices. We extend this idea to MED by reparametrizing the scale matrices ${\boldsymbol{\Sigma}}_{k}$ as follows,
@@
where ${\mathb{D}}_{k}$ is a $M\times M$ orthogonal matrix which contains the eigenvectors of ${\boldsymbol{\Sigma}}_{k}$ and ${\mathb{A}}_{k}$ is a $M\times M$ diagonal matrix that contains the associated eigenvalues in decreasing order. The key idea introduced by Bouveyron, Girard and Schmid (2007) is to consider that each cluster lies in a low-dimensional subspace of dimension $d_{k}<M$, which can be expressed by assuming that
@@
where $a_{k1},\ldots,a_{kd_{k}}$ are the $d_{k}$ largest eigenvalues of ${\boldsymbol{\Sigma}}_{k}$ and $b_{k}$ is a small negligible value. The $d_{k}$ eigenvectors associated to the first $d_{k}$ eigenvalues $\{a_{k1},\ldots,a_{kd_{k}}\}$ define a cluster-specific subspace $\mathbb{E}_{k}$, which captures the main cluster shape. The orthogonal subspace is denoted by $\mathbb{E}_{k}^{\bot}$. Let $\widetilde{{\mathb{D}}}_{k}$ consists of the $d_{k}$ first columns of ${\mathb{D}}_{k}$ supplemented by $(M-d_{k})$ zero columns and $\overline{{\mathb{D}}}_{k}=({\mathb{D}}_{k}-\widetilde{{\mathb{D}}}_{k})$. It follows that $P_{k}(\mathbf{y})=\widetilde{{\mathb{D}}}_{k}\widetilde{{\mathb{D}}}_{k}^{T}(\mathbf{y}-{\boldsymbol{\mu}}_{k})+{\boldsymbol{\mu}}_{k}$ and $P_{k}^{\bot}(\mathbf{y})=\overline{{\mathb{D}}}_{k}\overline{{\mathb{D}}}_{k}^{T}(\mathbf{y}-{\boldsymbol{\mu}}_{k})+{\boldsymbol{\mu}}_{k}$ are the projections of $\mathbf{y}$ on $\mathbb{E}_{k}$ and $\mathbb{E}_{k}^{\bot}$ respectively. This parameterization allows to handle high dimensional data in a computationally efficient way. For instance the quadratic form of the Mahalanobis distance, appearing in the generator $g$ in (2), writes
@@
where .$\|.\|^{2}_{\widetilde{{\boldsymbol{\Sigma}}}_{k}^{-1}}$ is the norm defined by $\|\mathbf{y}\|^{2}_{\widetilde{{\boldsymbol{\Sigma}}}^{-1}_{k}}=\mathbf{y}^{T}\widetilde{{\boldsymbol{\Sigma}}}^{-1}_{k}\mathbf{y}$ with $\widetilde{{\boldsymbol{\Sigma}}}_{k}^{-1}=\widetilde{{\mathb{D}}}_{k}{\mathb{A}}_{k}^{-1}\widetilde{{\mathb{D}}}_{k}^{T}$. Equation (7) uses the definitions of $P_{k}$ and $P_{k}^{\bot}$ and $\|{\boldsymbol{\mu}}_{k}-P_{k}^{\bot}(\mathbf{y})\|^{2}=\|\mathbf{y}-P_{k}(\mathbf{y})\|^{2}$. The gain comes from the fact that (7) does not depend on $P_{k}^{\bot}$ and thus does not require the computation of the $(M-d_{k})$ latest columns of ${\mathb{D}}_{k}$, the eigenvectors associated to the smallest eigenvalues. Similarly, determinants can be efficiently computed as $\text{log}(|{\boldsymbol{\Sigma}}_{k}|)=(\sum_{m=1}^{d_{k}}\text{log}(a_{km}))+(M-d)\text{log}(b_{k})$. This efficient parameterization, that now only depends on matrix $\widetilde{{\mathb{D}}}_{k}$ and not on the complete matrix ${\mathb{D}}_{k}$, is indicated by denoting the corresponding ED as $\mathcal{HE}_{Md_{k}}\left({\boldsymbol{\mu}}_{k},\widetilde{{\mathb{D}}}^{*}_{k},{\mathb{a}}_{k},b_{k},{\boldsymbol{\theta}}_{k}\right)$, where $\widetilde{{\mathb{D}}}^{*}_{k}$ the matrix $\widetilde{{\mathb{D}}}_{k}$ with the last zeros $M-d_{k}$ columns omitted. We then refer to a MED coupled with this parameterization as a high-dimensional MED (HD-MED).
@@
A random vector $\mathbf{Y}\in\mathbb{R}^{M}$ follows a HD-MED distribution if for all $k\in[1:K]$, the conditional pdf of the $k^{\text{th}}$ mixture component $f_{k}$ is an ED $\mathcal{HE}_{Md_{k}}\left({\boldsymbol{\mu}}_{k},\widetilde{{\mathb{D}}}^{*}_{k},{\mathb{a}}_{k},b_{k},{\boldsymbol{\theta}}_{k}\right)$ with reparameterization given by (5) and (6). We denote
$\mathbf{Y}\sim\mathcal{MHE}_{M{\mathb{d}}}\left((\pi_{k},{\boldsymbol{\mu}}_{k},\widetilde{{\mathb{D}}}^{*}_{k},{\mathb{a}}_{k},b_{k},{\boldsymbol{\theta}}_{k})_{k=1}^{K}\right).$
With ${\mathb{d}}=\left(d_{1},\dots,d_{k}\right)$, ${\mathb{a}}_{k}=\left(a_{k1},\dots,a_{kd_{k}}\right)$.
@@ 5.1 Latent variable dimension reduction
Standard PCA is defined without referring to a probabilistic model. Given a set of observations in $\mathbb{R}^{M}$, their $M\times M$ empirical covariance matrix is decomposed into eigenvalues and eigenvectors and a number $d<<M$ of them are retained. For any observation $\mathbf{y}\in\mathbb{R}^{M}$, a lower dimensional representation can then be obtained by considering its projection to a lower dimensional subspace $\widehat{\mathbf{y}}=\hat{{\boldsymbol{\Sigma}}}_{d}^{T}\mathbf{y}$ where $\widehat{{\boldsymbol{\Sigma}}}_{d}$ is the matrix containing the $d$ first eigenvectors of the empirical covariance matrix. If needed, its reconstruction in $\mathbb{R}^{M}$, optimal in the sense of the squared reconstruction error, can be obtained with $\widetilde{\mathbf{y}}=\widehat{{\boldsymbol{\Sigma}}}_{d}\widehat{\mathbf{y}}$. Alternatively, if $\mathbf{y}$ is assumed to be a realization of a random vector $\mathbf{Y}\sim\mathcal{HE}_{Md}\left({\boldsymbol{\mu}},\widetilde{{\mathb{D}}}^{*},{\mathb{a}},b,{\boldsymbol{\theta}}\right)$, a low dimensional representation of $\mathbf{y}$ can be justified using the following latent variable model representation of $\mathbf{Y}$.
@@
Let $d\leq M-1$, $\mathbf{Y}\in\mathbb{R}^{M}$, $\mathbf{X}\in\mathbb{R}^{d}$, ${\mathb{E}}\in\mathbb{R}^{M}$, $W\in\mathbb{R}^{+}$ be random variables, ${\mathb{V}}\in\mathbb{R}^{M\times d}$ a matrix of linearly independent columns, ${\boldsymbol{\mu}}\in\mathbb{R}^{M}$ a vector and $f_{\boldsymbol{\theta}}$ the pdf of a positive univariate random variable defined by some parameter ${\boldsymbol{\theta}}$. Assume that
$\mathbf{Y}$ $={\mathb{V}}\mathbf{X}+{\boldsymbol{\mu}}+{\mathb{E}}$
$(\mathbf{X}|W=w)$ $\sim{\cal N}(\mathbf{0}_{d},w^{-1}{\mathb{I}}_{d})$
$({\mathb{E}}|W=w)$ $\sim{\cal N}(\mathbf{0}_{M},b\;w^{-1}{\mathb{I}}_{M})$
$W$ $\sim f_{\boldsymbol{\theta}}$
then,
$\mathbf{Y}\sim\mathcal{HE}_{Md}\left({\boldsymbol{\mu}},\widetilde{{\mathb{D}}}^{*},{\mathb{a}},b,{\boldsymbol{\theta}}\right),$
with $b{\mathb{I}}_{M}+{\mathb{V}}{\mathb{V}}^{T}={\mathb{D}}{\mathb{A}}{\mathb{D}}^{T}$ and ${\mathb{A}}=\operatorname{diag}(a_{1},\dots,a_{d},b,\dots,b)$.
@@
and from the GSM definition 4.2 that
$\mathbf{Y}\sim{\cal E}_{M}({\boldsymbol{\mu}},{\mathb{V}}{\mathb{V}}^{T}+b{\mathb{I}}_{M},f_{\boldsymbol{\theta}})\;.$
Since ${\mathb{V}}{\mathb{V}}^{T}$ is of rank $d$, ${\mathb{V}}{\mathb{V}}^{T}+b{\mathb{I}}_{M}$ admits an eigenvalue decomposition ${\mathb{D}}{\mathb{A}}{\mathb{D}}^{T}$ as stated.
Distribution (8) follows from standard Gaussian vectors properties. Using the tower property, it comes then $\mathbb{E}[\mathbf{Y}|\mathbf{X}=\mathbf{x}]=\mathbb{E}[\mathbb{E}[\mathbf{Y}|\mathbf{X}=\mathbf{x},W]]={\mathb{V}}\mathbf{x}+{\boldsymbol{\mu}}$ and $\mathbb{E}[\mathbf{X}|\mathbf{Y}=\mathbf{y}]=\mathbb{E}[\mathbb{E}[\mathbf{X}|\mathbf{Y}=\mathbf{y},W]]={\mathb{U}}^{-1}{\mathb{V}}^{T}(\mathbf{y}-{\boldsymbol{\mu}}).$
The previous proposition states that a realization $\mathbf{y}\in\mathbb{R}^{M}$ from an HD-ED, can also be seen as originating from a generative model with a lower-dimensional latent variable $\mathbf{X}\in\mathbb{R}^{d}$. Hence, a natural alternative to the standard PCA projection, is the conditional mean $\widehat{\mathbf{y}}=\mathbb{E}[\mathbf{X}|\mathbf{Y}=\mathbf{y}]$, that is
@@
and as a reconstruction or an approximation of the original information, the conditional mean $\widetilde{\mathbf{y}}=\mathbb{E}[\mathbf{Y}|\mathbf{X}=\widehat{\mathbf{y}}]$, that is $\widetilde{\mathbf{y}}={\mathb{V}}\widehat{\mathbf{y}}+{\boldsymbol{\mu}}$. Using the previous formulas, lower dimensional representations can thus be obtained using ${\mathb{U}}$ and ${\mathb{V}}$ but when estimating the parameters of the HD-ED, we get estimates for ${\mathb{A}}$ and ${\mathb{D}}$ instead. However, using that ${\mathb{V}}$ is of rank $d$ and ${\mathb{D}}{\mathb{A}}{\mathb{D}}^{T}=b{\mathb{I}}_{M}+{\mathb{V}}{\mathb{V}}^{T}$, we can set
@@
and deduce ${\mathb{U}}$ straightforwardly.

### v2

@@ 3.2 Contribution
Herein, we propose to explore this divide-and-conquer strategy by introducing the framework of High Dimensional Mixtures of Elliptical Distributions (HD-MED) to overcome the challenge of estimating more than $2$ parameters, e.g. $T_{1}$, $T_{2}$, $\delta\!f$, $B_{1}$, CBV, $R$. Considering the reconstruction of 6 parameters makes the associated dictionaries both large in size (an order of $13$ terabytes of signals to represent the tissue heterogeneity) and dimension, framing our proposal around two key components. Building on HD-MED, a generalization of HD-GMM, we show how they can be used to simultaneously compress and cluster large-scale high-dimensional MRF dictionaries, and more generally any dataset. We derive a new online algorithm, based on a principled EM framework, to learn such a model from very large data volumes. We demonstrate the effectiveness of our approach on MRF reconstruction, showing results comparable to the high-dimensional dictionary matching referred to as full-matching in the next sections. This approach allows us to exceed the resolution, size used in current implementations and to reconstruct a larger number of MR parameter maps with improved accuracy, thereby advancing the clinical feasibility of MRF.
@@ 4 High Dimensional Mixtures of Elliptical Distributions
Due to time constraints and computational complexity, data may be acquired with errors and at high noise levels. In such cases, using Gaussian models, which are known to be sensitive to outliers, is not recommended. Elliptical distributions represent a family of distributions that contains Gaussian distributions but also heavy-tailed distributions, such as the Student or Laplace distributions. Elliptical distributions allow more flexible modeling, with additional parameters and better fit to data than Gaussian distributions, see e.g. (Lange1989; Lange1993; Neri2021; Kotz2001). This additional flexibility is often referred to as robustness because the heavier tails make estimation less sensitive to outlying samples. In this paper, we consider a sub-class of elliptical distributions, which can be expressed as infinite mixtures of Gaussian distributions.
@@ 4.1 Gaussian scale mixture (GSM) distributions
Scale mixtures of multivariate Gaussian distributions, also referred to as normal/independent distributions (Lange1993), are an important subclass of elliptical distributions whose definition is recalled below. Scale mixtures of Gaussians share good properties with Gaussian distributions. They are tractable, lead to tractable inference procedures and provide more robust results, in contrast to Gaussian distributions that usually suffer from sensitivity to outliers.
@@
A continuous random vector $\mathbf{Y}\in\mathbb{R}^{M}$ follows a multivariate elliptical symmetric distribution if its probability density function (pdf) $p(\mathbf{y})$ is of the following form (see cambanis or kelker),
@@
Note that the scale matrix $\boldsymbol{\Sigma}$} is not necessarily the covariance matrix, $\boldsymbol{\Sigma}$} is proportional to the covariance matrix if the latter exists. The pdf of the multivariate normal distribution is a special case of ED with $g(u)=\exp(-u^{2}/2)$. Another member of the elliptical family is the multivariate Student distribution. This distribution is well-studied in the literature (kotz2004multivariate) and admits a useful representation as a Gaussian scale mixture. Denoting ${\cal N}_{M}(\mathbf{y};{\boldsymbol{\mu}},{\boldsymbol{\Sigma}})$ a $M$-variate Gaussian distribution with mean $\boldsymbol{\mu}$} and covariance matrix $\boldsymbol{\Sigma}$}, a Gaussian scale mixture distribution is a distribution of the following form.
@@
If $\boldsymbol{\mu}$} is an $M$-dimensional vector, $\boldsymbol{\Sigma}$} is an $M\times M$ positive definite symmetric matrix and $f$ is a pdf of a univariate positive variable $W\in\mathbb{R}^{+}$, then the $M$-dimensional density given by
$p(\mathbf{y})=\int_{\mathbb{R}^{+}}{\cal N}_{M}\left(\mathbf{y};{\boldsymbol{\mu}},\frac{{\boldsymbol{\Sigma}}}{w}\right)\;f(w)\;dw$ (3)
@@
In practice, we will consider mixing distribution $f_{\boldsymbol{\theta}}$ that depends on some parameter $\boldsymbol{\theta}$} and also write $\mathbf{Y}\sim{\cal E}_{M}\left({\boldsymbol{\mu}},{\boldsymbol{\Sigma}},{\boldsymbol{\theta}}\right)$. As already mentioned, famous GSM distributions include the multivariate Student distribution (when $f_{\boldsymbol{\theta}}$ is the pdf of a chi-squared variable), the Pearson type VII distribution (when $f_{\boldsymbol{\theta}}$ is a gamma distribution), the generalized Gaussian (when a power of $W$ follows the gamma distribution) and the Laplace distribution (when $W^{-1}$ follows an exponential distribution). It is straightforward to see that GSM are elliptical distributions. However, not all elliptical distributions can be reduced to scale mixtures. For the previous reason, characterization and a way to represent elliptical distributions as GSM are very valuable. In Gomez2006, conditions are given under which elliptical distributions are GSMs. The issue of finding the corresponding mixing distribution is also addressed. An illustration of these results for generalized Gaussian distributions is given by Gomez2008.
@@ 4.2 Mixtures of High Dimensional GSM
We consider finite mixtures of $K$ GSM distributions, with a specific parameterization to handle potentially high-dimensional observations.
Definition 4.3 (Finite mixture).
@@
with $\pi_{k}\in[0,1]$ the mixing weights summing to one and $f_{k}$ the pdf of the $k^{\text{th}}$ component.
A mixture model where each $f_{k}$ is a GSM (3), ${\cal E}({\boldsymbol{\mu}}_{k},{\boldsymbol{\Sigma}}_{k},{\boldsymbol{\theta}}_{k})$, is referred to as a mixture of ED (MED). The number of parameters in MED grows quadratically with the dimension $M$ due to the scale matrices ${\boldsymbol{\Sigma}}_{k}$. For large $M$, this can be problematic for estimation. In the Gaussian case, to reduce the number of parameters, bouveyron2007 proposed a family of parsimonious Gaussian mixtures, using the eigenvalue decomposition of the covariance matrices. We extend this idea to MED by reparameterizing ${\boldsymbol{\Sigma}}_{k}$ as follows,
@@
where ${\mathb{D}}_{k}$ is an $M\times M$ orthogonal matrix, which contains the eigenvectors of ${\boldsymbol{\Sigma}}_{k}$ and ${\mathb{A}}_{k}$ is an $M\times M$ diagonal matrix, which contains the associated eigenvalues in decreasing order. The key idea of bouveyron2007 is to consider that each cluster lies in a low-dimensional subspace of dimension $d_{k}<M$, which can be expressed by assuming that
@@
where $a_{k1},\ldots,a_{kd_{k}}$ are the $d_{k}$ largest eigenvalues of ${\boldsymbol{\Sigma}}_{k}$ and $b_{k}$ is a small negligible value. The $d_{k}$ eigenvectors associated to the first $d_{k}$ eigenvalues $\{a_{k1},\ldots,a_{kd_{k}}\}$ define a cluster-specific subspace $\mathbb{E}_{k}$, which captures the main cluster shape. The orthogonal subspace is denoted by $\mathbb{E}_{k}^{\bot}$. Let $\widetilde{{\mathb{D}}}_{k}$ consists of the $d_{k}$ first columns of ${\mathb{D}}_{k}$ supplemented by $(M-d_{k})$ zero columns and $\overline{{\mathb{D}}}_{k}=({\mathb{D}}_{k}-\widetilde{{\mathb{D}}}_{k})$. It follows that $P_{k}(\mathbf{y})=\widetilde{{\mathb{D}}}_{k}\widetilde{{\mathb{D}}}_{k}^{T}(\mathbf{y}-{\boldsymbol{\mu}}_{k})+{\boldsymbol{\mu}}_{k}$ and $P_{k}^{\bot}(\mathbf{y})=\overline{{\mathb{D}}}_{k}\overline{{\mathb{D}}}_{k}^{T}(\mathbf{y}-{\boldsymbol{\mu}}_{k})+{\boldsymbol{\mu}}_{k}$ are the projections of $\mathbf{y}$ on $\mathbb{E}_{k}$ and $\mathbb{E}_{k}^{\bot}$ respectively. This parameterization allows us to handle high-dimensions in an efficient way. For instance, the quadratic form of the Mahalanobis distance, appearing in the generator $g$ in (2), writes
@@
where $\|.\|^{2}_{\widetilde{{\boldsymbol{\Sigma}}}_{k}^{-1}}$ is the norm defined by $\|\mathbf{y}\|^{2}_{\widetilde{{\boldsymbol{\Sigma}}}^{-1}_{k}}=\mathbf{y}^{T}\widetilde{{\boldsymbol{\Sigma}}}^{-1}_{k}\mathbf{y}$ with $\widetilde{{\boldsymbol{\Sigma}}}_{k}^{-1}=\widetilde{{\mathb{D}}}_{k}{\mathb{A}}_{k}^{-1}\widetilde{{\mathb{D}}}_{k}^{T}$. Equation (7) uses the definitions of $P_{k}$ and $P_{k}^{\bot}$ and $\|{\boldsymbol{\mu}}_{k}-P_{k}^{\bot}(\mathbf{y})\|^{2}=\|\mathbf{y}-P_{k}(\mathbf{y})\|^{2}$. The gain comes from the fact that (7) does not depend on $P_{k}^{\bot}$ and thus does not require the computation of the $(M-d_{k})$ latest columns of ${\mathb{D}}_{k}$, the eigenvectors associated to the smallest eigenvalues. Similarly, determinants can be efficiently computed as $\text{log}(|{\boldsymbol{\Sigma}}_{k}|)=(\sum_{m=1}^{d_{k}}\text{log}(a_{km}))+(M-d)\text{log}(b_{k})$. This parameterization that now only depends on $\widetilde{{\mathb{D}}}_{k}$ and not on the complete ${\mathb{D}}_{k}$, is indicated by denoting the corresponding ED as $\mathcal{HE}_{Md_{k}}\left({\boldsymbol{\mu}}_{k},\widetilde{{\mathb{D}}}^{*}_{k},{\mathb{a}}_{k},b_{k},{\boldsymbol{\theta}}_{k}\right)$, where $\widetilde{{\mathb{D}}}^{*}_{k}$ is matrix $\widetilde{{\mathb{D}}}_{k}$ with the last zero $M-d_{k}$ columns omitted. We then refer to a MED coupled with this parameterization as a high-dimensional MED (HD-MED).
@@
A random vector $\mathbf{Y}\in\mathbb{R}^{M}$ follows a HD-MED distribution if for all $k\in[1:K]$, the pdf of the $k^{\text{th}}$ component $f_{k}$ is an ED $\mathcal{HE}_{\!Md_{k}}\!\!\left({\boldsymbol{\mu}}_{k},\widetilde{{\mathb{D}}}^{*}_{k},{\mathb{a}}_{k},b_{k},{\boldsymbol{\theta}}_{k}\right)$ with reparameterization given by (5) and (6). We denote
$\mathbf{Y}\sim\mathcal{MHE}_{M{\mathb{d}}}\left((\pi_{k},{\boldsymbol{\mu}}_{k},\widetilde{{\mathb{D}}}^{*}_{k},{\mathb{a}}_{k},b_{k},{\boldsymbol{\theta}}_{k})_{k=1}^{K}\right),$ with ${\mathb{d}}=\left(d_{1},\dots,d_{k}\right)$, ${\mathb{a}}_{k}=\left(a_{k1},\dots,a_{kd_{k}}\right)$.
@@ 5.1 Latent variable dimension reduction
Standard PCA is defined without referring to a probabilistic model. Given a set of observations in $\mathbb{R}^{M}$, their $M\times M$ empirical covariance matrix is decomposed into eigenvalues and eigenvectors. For any $\mathbf{y}\in\mathbb{R}^{M}$, a lower dimensional ($d<<M$) representation can then be obtained by considering its projection to a lower dimensional subspace $\widehat{\mathbf{y}}=\hat{{\boldsymbol{\Sigma}}}_{d}^{T}\mathbf{y}$ where $\widehat{{\boldsymbol{\Sigma}}}_{d}$ is the matrix containing the $d$ first eigenvectors of the empirical covariance matrix. Its reconstruction in $\mathbb{R}^{M}$, optimal in the sense of the squared reconstruction error, can be obtained with $\widetilde{\mathbf{y}}=\widehat{{\boldsymbol{\Sigma}}}_{d}\widehat{\mathbf{y}}$. Alternatively, if $\mathbf{y}$ is assumed to be a realization of a random vector $\mathbf{Y}\sim\mathcal{HE}_{Md}\left({\boldsymbol{\mu}},\widetilde{{\mathb{D}}}^{*},{\mathb{a}},b,{\boldsymbol{\theta}}\right)$, a low dimensional representation of $\mathbf{y}$ can be justified using the following latent variable model representation of $\mathbf{Y}$. The following proposition generalizes the developments of archambeau2006; Guo2023 for the Student distribution.
@@
Let $d\leq M-1$, $\mathbf{Y}\in\mathbb{R}^{M}$, $\mathbf{X}\in\mathbb{R}^{d}$, ${\mathb{E}}\in\mathbb{R}^{M}$, $W\in\mathbb{R}^{+}$ be random variables, ${\mathb{V}}\in\mathbb{R}^{M\times d}$ a matrix of linearly independent columns, ${\boldsymbol{\mu}}\in\mathbb{R}^{M}$ a vector and $f_{\boldsymbol{\theta}}$ the pdf of a positive univariate random variable defined by some parameter $\boldsymbol{\theta}$}. Assume that $\mathbf{Y}={\mathb{V}}\mathbf{X}+{\boldsymbol{\mu}}+{\mathb{E}}$, $(\mathbf{X}|W=w)\sim{\cal N}(\mathbf{0}_{d},w^{-1}{\mathb{I}}_{d})$, $({\mathb{E}}|W=w)\sim{\cal N}(\mathbf{0}_{M},b\;w^{-1}{\mathb{I}}_{M})$ and $W\sim f_{\boldsymbol{\theta}}$, then, $\mathbf{Y}\sim\mathcal{HE}_{Md}\left({\boldsymbol{\mu}},\widetilde{{\mathb{D}}}^{*},{\mathb{a}},b,{\boldsymbol{\theta}}\right),$ with ${\mathb{D}}{\mathb{A}}{\mathb{D}}^{T}=b{\mathb{I}}_{M}+{\mathb{V}}{\mathb{V}}^{T}$ the eigenvalue decomposition of $b{\mathb{I}}_{M}+{\mathb{V}}{\mathb{V}}^{T}$, ${\mathb{A}}=\operatorname{diag}(a_{1},\dots,a_{d},b,\dots,b)$ the ordered eigenvalues and $\widetilde{{\mathb{D}}}^{*}$ the matrix containing the first $d$ eigenvectors as columns.
@@
and from definition 4.2 that $\mathbf{Y}\sim{\cal E}_{M}({\boldsymbol{\mu}},{\mathb{V}}{\mathb{V}}^{T}+b{\mathb{I}}_{M},f_{\boldsymbol{\theta}})\;.$ Since ${\mathb{V}}{\mathb{V}}^{T}$ is of rank $d$, ${\mathb{V}}{\mathb{V}}^{T}+b{\mathb{I}}_{M}$ admits an eigenvalue decomposition ${\mathb{D}}{\mathb{A}}{\mathb{D}}^{T}$ as stated. Distribution (8) follows from standard Gaussian vectors properties. Using the tower property, $\mathbb{E}[\mathbf{Y}|\mathbf{X}=\mathbf{x}]=\mathbb{E}[\mathbb{E}[\mathbf{Y}|\mathbf{X}=\mathbf{x},W]]={\mathb{V}}\mathbf{x}+{\boldsymbol{\mu}}$ and $\mathbb{E}[\mathbf{X}|\mathbf{Y}=\mathbf{y}]=\mathbb{E}[\mathbb{E}[\mathbf{X}|\mathbf{Y}=\mathbf{y},W]]={\mathb{U}}^{-1}{\mathb{V}}^{T}(\mathbf{y}-{\boldsymbol{\mu}}).$
The previous proposition states that $\mathbf{y}\in\mathbb{R}^{M}$ from an HD-ED, can also be seen as originating from a model with a lower-dimensional latent variable $\mathbf{X}\in\mathbb{R}^{d}$. Hence, a natural alternative to the standard PCA projection, is the conditional mean $\widehat{\mathbf{y}}=\mathbb{E}[\mathbf{X}|\mathbf{Y}=\mathbf{y}]$, that is
@@
and as a reconstruction or an approximation of the original information, the conditional mean $\check{\mathbf{y}}=\mathbb{E}[\mathbf{Y}|\mathbf{X}=\widehat{\mathbf{y}}$], that is $\check{\mathbf{y}}={\mathb{V}}\widehat{\mathbf{y}}+{\boldsymbol{\mu}}$. Using the previous formulas, lower dimensional representations can thus be obtained using $\mathb{U}$} and $\mathb{V}$} but when estimating the parameters of the HD-ED, we get estimates for $(a_{1},\dots,a_{d})$, $b$ and $\widetilde{{\mathb{D}}}^{*}$. However, using that $\mathb{V}$} is of rank $d$ and ${\mathb{D}}{\mathb{A}}{\mathb{D}}^{T}=b{\mathb{I}}_{M}+{\mathb{V}}{\mathb{V}}^{T}$, we can set
@@
and deduce $\mathb{U}$} straightforwardly. Interestingly, the projection and reconstruction formulas do not directly depend on $f_{\boldsymbol{\theta}}$. The dependence is only implicit via the impact of $f_{\boldsymbol{\theta}}$ on the estimation of the other parameters.